OpenStudy (anonymous):

How to find the average rate of change from f(x)=-x^2+4x+1 from x=3 to x=6?

jimthompson5910 (jim_thompson5910):

are you able to compute the value of `f(3)` ?

jimthompson5910 (jim_thompson5910):

if not, then replace every x with 3 f(x)=-x^2+4x+1 f(3)=-(3)^2+4(3)+1 ... replace every x with 3 f(3) = _______ (fill in the blank and tell me what you get)

OpenStudy (anonymous):

it would equal 4

jimthompson5910 (jim_thompson5910):

correct, how about `f(6)` ?

OpenStudy (anonymous):

-11

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now to find the average rate of change, we use this formula \[\Large V = \frac{f(a)-f(b)}{a-b}\] V = average rate of change from x = a to x = b. I didn't want to use A because 'a' is already taken

jimthompson5910 (jim_thompson5910):

so, \[\Large V = \frac{f(a)-f(b)}{a-b}\] \[\Large V = \frac{f(3)-f(6)}{3-6}\] \[\Large V = \frac{4 - (-11)}{3-6}\] I'll let you finish up

OpenStudy (anonymous):

so v=-5, right?

jimthompson5910 (jim_thompson5910):

correct, the average rate of change from x = 3 to x = 6 is `-5` in a visual sense, this means that the slope of the line through the two points A and B on f(x) is going to be -5 A is the point corresponding to x = 3 B is the point corresponding to x = 6